OpenStudy (anonymous):

4 sin(x) sin(pi/3+x) sin(pi/3-x) = sin(3 x)

OpenStudy (anonymous):

2sinx(2sinAsinB) => 2sinx(cos(A-B) - cos(A+B)) A-B = pi/3 + x - pi/3 + x = 2x A+B = pi/3 +x + pi/3 - x = 2pi/3 => cos(pi - pi/3) => -cos pi/3 => -1/2 2sinx(cos2x + 1/2) => 2sinxcos2x + sinx => 2sinx(2cos^2x - 1) + sinx => 4sinxcos^2x - sinx => 4sinx(1-sin^2x) - sinx => 3sinx - 4sin^3x =>sin3x Hope You Understood

OpenStudy (anonymous):

thank you i have been stuck on that one forever.

OpenStudy (anonymous):

Haha... You're Welcome :)

OpenStudy (anonymous):

sorry to bother you again but what identity formulas did you use.

OpenStudy (anonymous):

2sinAsinB = cos(A-B) - cos(A+B) cos2x = 2cos^2x - 1 = 1- 2sin^2x cos^2x = 1 - sin^2x sin3x = 3sinx - 4sin^3x These are the things I used.

OpenStudy (anonymous):

thanks